# Unisolver: PDE-Conditional Transformers Are Universal PDE Solvers

### v3

@@ Results
4.2 1D Time-Dependent PDEs
Figure 4: Showcases from various evaluation scenarios on the 1D time-dependent PDEs.
Setups
This benchmark contains three million high-quality 1D time-dependent PDE samples with varying equation formulations, coefficients, force terms and boundary conditions. We perform joint training on this extensive dataset, where the input for the training task includes all relevant PDE components, and the output records full space-time fields. After training, the model is evaluated across multiple test settings, including in-distribution test, as well as zero-shot generalization on the Burgers and Advection equations from PDEBench (Takamoto et al., 2022), which is an another unseen dataset.
1D Time-dep endent PDEs Params Tasks In-distribution Test
Zero-shot Burgers
Zero-shot Advection
$\nu=0.1$ $\nu=0.01$ $\nu=0.001$ $\beta=0.1$
PDEformer 22M 0.0225 0.00744 0.0144 0.0393 0.0178
Unisolver 19M 0.0108 0.00513 0.00995 0.0299 0.0138
Promotion / 52.0% 31.0% 30.9% 23.9% 22.5%
Results
Table 5 presents that the in-distribution test performance of Unisolver is significantly better than that of PDEformer, indicating that our design of incorporating complete PDE components is more effective than the computational graph utilized by PDEformer in representing intricate physical relations. Additionally, Unisolver achieves better performance in four zero-shot generalization scenarios, with an average improvement of 27.1% over PDEformer, even with fewer parameters.
4.3 2D Mixed PDEs
Figure 5: Showcases from in-distrib-ution test sets on the 2D mixed PDEs.
Setups
This benchmark involves 12 datasets from four prominent benchmarks, covering a wide range of PDEs. After joint training on these diverse datasets, we perform in-distribution tests on each dataset. Notably, the in-distribution test set also involves challenging variations in the PDE components. Moreover, unlike the balanced data in HeterNS, these datasets exhibit significant imbalances across different PDE components. To mitigate this issue, we adopt the balanced data sampling method from DPOT (Hao et al., 2024); however, it still poses considerable challenges in managing such diverse PDE samples.
Results
As shown in Table 6, Unisolver outperforms DPOT (Hao et al., 2024) in 11 out of 12 in-distribution test sets with an remarkable average promotion of 17.5% (5.50→ 4.54), except for the small Diffusion-Reaction (DR) dataset whose relative L2 is less than 5%, verifying the effectiveness of our design in modeling such complex relations. Unisolver shows consistently superior performance in PDE component-dominated tasks, including coefficient generalization in FNO (2021a), force generalization in PDEArena (2023), and geometry generalization in CFDBench (2023), highlighting its ability to capture generalizable representations from complete PDE components.
Table 6: Performance comparison (relative L2 (×$\times 10^{-2}$)) across 12 in-distribution test sets. For conciseness, we use a “source-PDE” format to denote different tasks (e.g. FNO-NS). The second row lists the primary PDE components considered for each dataset. See Appendix E.3 for details.
2D Mixed PDEs Params Tasks
FNO-NS-$\nu$
PDEBench-CNS-(M,$\zeta$)
PDEBench
PDEArena
CFDBench-NS
Average Error
1e-5
1e-4
1e-3
(1, 0.1)
(1, 0.01)
(0.1, 0.1)
(0.1, 0.01)
DR
SWE
NS
NS-Force
Geometry
DPOT
30M
5.53
4.42
1.31
1.53
3.37
1.19
1.87
3.79
0.66
9.91
31.6
0.70
5.50
Unisolver
33M
4.17
3.36
0.61
1.23
2.89
1.01
1.59
4.39
0.45
6.87
27.4
0.54
4.54
Promotion (%)
24.6
24.0
53.4
19.6
14.2
15.1
15.0
31.8
30.7
13.3
22.9
17.5
4.4 Model Analysis
Ablations
As shown in Table 7, we further investigate the effect of LLM embeddings and condition modeling modules on 50,000 samples from the 1D time-dependent PDEs benchmark.
Firstly, in the LLM ablations, without LLM embedding, performance is the worst among all cases, even worse than replacing by random vector. LLaMA-3 brings a remarkable 5.76% averaged promotion compared to models without LLM embedding, indicating its essential role in learning PDEs. Notably, since the LLM only encodes one of six components, the equation symbols, a promotion of around 5% is a significant margin. Moreover, we compare the Unisolver’s performance across different language models in Figure 6, including LLaMA-3, LLaMA-2 and T5. The results are comparable, indicating each model possesses sufficient ability to encode prior mathematical knowledge.
Secondly, in condition modeling ablations, removing subspace decoupling introduces interference between different groups of PDE conditions, significantly impairing performance in zero-shot generalization settings, with an average drop of 5.45%. Moreover, direct concatenation of PDE components severely hinders relation learning (21.0% average drop), indicating the benefits of our design.
Table 7: Ablation results on the LLM embeddings and the Condition Modeling. Variants of the former include without LLM embeddings (w/o LLM) and replacing by random vectors (Random Vector), and variants of the latter include without subspace decoupling (w/o Subspace) and directly concatenating components (Concat). “Unchanged” means no changes to the default design.
1D Time- dependent PDEs LLM Embeddings Condition Modeling In-distribution Test Zero-shot Generalization
Burgers Burgers Burgers Advection
$\nu=0.1$ $\nu=0.01$ $\nu=0.001$ $\beta=0.1$
Random Vector Unchanged 0.0290 0.0185 0.0675 0.1471 0.0632
Unchanged w/o Subspace 0.0287 0.0187 0.0675 0.1478 0.0625
Unchanged Concat 0.0317 0.0236 0.0802 0.1586 0.0732
final Unchanged Unchanged 0.0277 0.0176 0.0659 0.1350 0.0603
Figure 6: (a) Comparison of different language models. (b) PCA visualization of LLM embeddings. The considered PDE family contains six coefficients, such as $c_{01},c_{02},c_{03}$. Different colors represent the number of non-zero coefficients, intuitively indicating the complexity of PDEs. A zero coefficient results in the removal of a term from the equation, impacting the representations embedded by the LLM. (c) PCA visualization of learned deep PDE conditions, $\mathbf{I}_{\text{select}}$ in Eq. (3). We vary only one coefficient at a time and keep the others fixed at zero, forming the shown parabolic-like trajectories.
Visualization of the learned PDE embeddings
As depicted in Figure 6(b-c), we apply the principal component analysis (PCA) (Jolliffe & Cadima, 2016) to intuitively visualize the LLM embeddings of equation symbols and deep PDE conditions learned by Unisolver for 1D time-dependent PDEs. In Figure 6(b), we observe that PDEs with similar complexity are encoded into similar embeddings, highlighting that LLM can indeed effectively capture prior mathematical knowledge. In Figure 6(c), the trajectories of deep conditions resemble parabolas with varying degrees of curvature, indicating that the learned deep conditions successfully capture the variations of PDE components.
Incomplete component scenario
In real-world applications, we may lack complete PDE components. To demonstrate the capability of Unisolver in handling such situations, we randomly replace PDE components with learnable tokens at a 30% probability in the HeterNS benchmark to simulate partially observed real-world data. For inference, we can flexibly choose whether to provide PDE components as inputs. As shown in Figure 7, even with incomplete components, Unisolver surpasses FNO (2021a) in most cases, especially in more complex cases with lower viscosity coefficients. Moreover, complete PDE information further improves the model’s performance (average 21.6%), supporting our motivation that complete information is essential for PDE solving.
Scalability
Figure 8: Data scalability (60x) and model scalability (21x) on the 1D time-dependent PDEs. Relative L2 results are plotted on a log-log scale.
Scalability is crucial for building a universal PDE solver. Figure 8 illustrates our exploration of Unisolver’s scalability, where we progressively increase the training data by 60 times and the model parameters by 21 times. Unisolver exactly displays the scaling law, achieving better performance with increased data and parameters, posing the potential for a practically universal PDE solver.
Case study
To provide a clear comparison, we provide showcases on the HeterNS in Figure 9. All the presented trajectories are generated from the same initial condition but exhibit distinct final fields, underscoring the determining role of PDE components. Further, we observe that Unisolver significantly outperforms FNO under complex conditions, such as smaller viscosity $\nu$ and larger force coefficient $\omega$, particularly in zero-shot generalization settings. More showcases can be found in Appendix B.
Figure 9: Error maps of FNO and Unisolver on the HeterNS, where all cases share the same initial condition but differ in viscosity ($\nu$) and force ($\omega$) (shown in the first row by the pairs ($\nu,\omega$)). The left panel shows in-distribution tests, while the right panel shows zero-shot generalization settings.
5 Conclusion
To break the generalization bottleneck, this paper presents Unisolver as a PDE-conditional Transformer, which stems from the theoretical analysis of the PDE-solving process. Concretely, Unisolver proposes a complete set of PDE components and embeds these components into domain-wise and point-wise deep conditions separately for deep conditions. By integrating these conditions with Transformers through a decoupled mechanism, Unisolver can handle universal PDE components and achieve consistent state-of-the-art results across three challenging, large-scale benchmarks. Extensive analyses are provided to verify the effectiveness, generalizability and scalability of our model.
Appendix A Fine-tuning Performance
Zero-shot generalization serves as a valuable metric, but in scenarios where datasets differ substantially from the training set, the model’s zero-shot performance may be limited. In such instances, fine-tuning performance is critical, since it reflects the model’s ability to learn fundamentally generalizable representations through large-scale training. We present Unisolver’s fine-tuning performance on 1D time-dependent and 2D mixed PDEs in Figures 10-11, with 100 epochs for 1D time-dependent PDEs and 200 epochs for 2D mixed PDEs, both amounting to 20% of the total training epochs from scratch, demonstrating fast adaptation.
Figure 10: Fine-tuning performance on 1D time-dependent PDEs. “FT-100” means fine-tuning on each dataset for 100 epochs.
For 1D time-dependent PDEs, as shown in Figure 10, fine-tuning 100 epochs on the Burgers and Advection equations from PDEBench (Takamoto et al., 2022) significantly enhances Unisolver’s performance, reducing error by 61% compared to zero-shot results and achieving a 59.3% improvement over PDEformer (Ye et al., 2024) under the same fine-tuning conditions. These results prove the condition modeling in Unisolver is more effective than the computational graph proposed by PDEformer, especially for fast adaptation. For 2D mixed PDEs, as shown in Figure 11, after 200 epochs of fine-tuning for each dataset, Unisolver reduces error by more than 12% compared to zero-shot generalization performance and outperforms DPOT (Hao et al., 2024) under the same fine-tuning conditions by 14%, showcasing its ability to extract generalizable knowledge from diverse training datasets.
Appendix B More Showcases
We provide additional showcases here to supplement the numerical results presented in the main text. First, we visualize the in-distribution test and zero-shot generalization cases on the HeterNS dataset in Figure 12 and Figure 13, respectively. Next, we present visualizations for 1D time-dependent PDEs in Figure 14. Finally, we illustrate the 12 diverse datasets from 2D mixed PDEs in Figure 15.
Figure 12: Showcases comparison (error maps) for in-distribution tests with top three baselines on the HeterNS dataset. See Table 3 and 4 for numerical comparison (relative L2). All data has the same initial condition and differs in viscosity ($\nu$) and force ($\omega$) (shown in the first row by the pairs ($\nu,\omega$)). Unisolver achieved the best visual performance among the compared baselines.
Figure 13: Showcases comparison (error maps) for zero-shot generalization settings with top three baselines on the HeterNS dataset with the same initial conditions and differs in viscosity ($\nu$) and force ($\omega$) (shown in the first row by the pairs ($\nu,\omega$)). See Table 3 and 4 for numerical comparison.
Figure 14: Error maps on the in-distribution test and zero-shot generalization (Burgers and Advection equation from PDEBench (Takamoto et al., 2022)) settings in 1D time-dependent PDEs. See Table 5 for numerical comparison. We visualize two cases: periodic boundary conditions and Robin boundary conditions in in-distribution tests. The number in the Burgers columns is the diffusion coefficient $\nu$ while the number in the Advection column is the advection speed $\beta$.
Figure 15: Unisolver predictions and error maps on 2D mixed PDEs. See Table 6 for numerical comparison with DPOT. Both predictions and error maps are provided. As shown in the CFDBench-NS columns, Unisolver presents an impressive ability to handle different geometry conditions.
Appendix C Full Trajectory Visualizations
To better understand the temporal evolution of the benchmark, we visualize the full trajectory of the ground truth and Unisolver predictions on HeterNS and 2D mixed PDEs. To further enhance clarity and provide a more intuitive understanding of these temporal dynamics, we have included videos to illustrate the trajectories frame by frame. Please refer to the supplementary material for details.
Appendix D Analytical Solution for the String Vibration Equation
We derive the explicit solution for Eq. (1a) with boundary conditions (1b) and initial conditions (1c).
We extend $\varphi(x)$, $\psi(x)$ and $f(x,t)$ as odd, periodic functions of $x$ with period $2L$, defining $\Phi(x)$, $\Psi(x)$ and $F(x,t)$ as follows:
$\Phi(x)=-\Phi(-x),\quad\Phi(x)=\Phi(x+2L),$ $-\infty<x<\infty,$ (4)
$\Phi(x)=\varphi(x),$ $0\leq x\leq L,$
$\Psi(x)=-\Psi(-x),\quad\Psi(x)=\Psi(x+2L),$ $-\infty<x<\infty,$
$\Psi(x)=\psi(x),$ $0\leq x\leq L.$
$F(x,t)=-F(-x,t),\quad F(x,t)=F(x+2L,t),$ $-\infty<x<\infty,0\leq t\leq T.$
$F(x,t)=f(x,t),$ $0\leq x\leq L,0\leq t\leq T.$
Thus, we obtain a wave equation defined on the upper half-plane, with initial conditions $\Phi(x)$ and $\Psi(x)$.
$\partial_{tt}U-a^{2}\partial_{xx}U=F,\quad(x,t)\in\mathbb{R}\times(0,T),$ (5)
$U(x,0)=\Phi(x),\quad\partial_{t}U(x,0)=\Psi(x),\quad x\in\mathbb{R}.$
Lemma 2.
The solution of the simplified equation (9) is $U(x,t)=\frac{1}{2a}\int_{x-at}^{x+at}\Psi(\xi)\mathrm{d}\xi$.
$\partial_{tt}U-a^{2}\partial_{xx}U=0,\quad(x,t)\in\mathbb{R}\times(0,T),$ (6)
$U(x,0)=0,\quad\partial_{t}U(x,0)=\Psi(x),\quad x\in\mathbb{R}.$
Proof.
The above expression for $U$ can be derived using the method of operator splitting and characteristic lines.
Note that
$0=\partial_{tt}U-c^{2}\partial_{xx}U=(\partial_{t}-a\partial_{x})(\partial_{t}+a\partial_{x})U.$
Let
$V=(\partial_{t}+a\partial_{x})U.$
Then
$\partial_{t}V-a\partial_{x}V=0,\quad V(x,0)=\Psi(x).$
Using the method of characteristic lines for first-order equations, we obtain
$V(x,t)=V(x+at,0)=\Psi(x+at).$
Thus, $U$ satisfies
$(\partial_{t}+a\partial_{x})U=V=\Psi(x+at).$
For this first-order equation, along the characteristic line passing through the point $(X,T^{\prime})$, defined by $x=at+X-aT^{\prime}$, we have
$\frac{\mathrm{d}u(X-aT^{\prime}+at,t)}{\mathrm{d}t}=\Psi(X-aT^{\prime}+2at).$
By integrating on both sides, we get
$U(X-aT^{\prime}+at,t)=\int_{0}^{t}\Psi(X-aT^{\prime}+2as)\,\mathrm{d}s.$
Therefore,
$U(x,t)=\int_{0}^{t}\Psi(x-at+2as)\,\mathrm{d}s=\frac{1}{2a}\int_{x-at}^{x+at}\Psi(\xi)\,\mathrm{d}\xi.$
It is straightforward to verify it’s indeed a solution by substituting it into the equation, given that $\Psi\in C^{1}(\mathbb{R})$. ∎
Inspired by the principle of linear superposition for linear problems, the solution $U=U(t,x)$ for Eq. (5) can be decomposed as
$U=U_{1}+U_{2}+U_{3},$
where $U_{1},U_{2},$ and $U_{3}$ are the solutions to the following three equations, respectively,
$\begin{cases}\partial_{tt}U-a^{2}\partial_{xx}U=0,\\ U(x,0)=\Phi(x),\\ \partial_{t}U(x,0)=0,\end{cases}$$\begin{cases}\partial_{tt}U-a^{2}\partial_{xx}U=0,\\ U(x,0)=0,\\ \partial_{t}U(x,0)=\Psi(x),\end{cases}$$\begin{cases}\partial_{tt}U-a^{2}\partial_{xx}U=F(x,t),\\ U(x,0)=0,\\ \partial_{t}U(x,0)=0.\end{cases}$
Lemma 3.
The three solutions have the following simple relationships. Denoting that $U_{2}=S_{\Psi}$, we have
[…]
Proof.
These relations can be verified by substitution. ∎
Thus, the explicit expression of $U$ is,
$U(x,t)$ $=(U_{1}+U_{2}+U_{3})(x,t)$
$=\frac{1}{2}(\Phi(x+at)+\Phi(x-at))+\frac{1}{2a}\int^{x+at}_{x-at}\Psi(\xi)\mathrm{d}\xi+\frac{1}{2a}\int_{0}^{t}\mathrm{d}\tau\int_{x-a(t-\tau)}^{x+a(t-\tau)}F(\xi,\tau)\mathrm{d}\xi.$
Noting that due to the construction of $\Phi(x)$ and $\Psi(x)$, the solution $U(x,t)$ of Eq. (5) is an odd function of $x$ with a period of $2L$. Moreover, it satisfies $U(kL,t)=0,\forall k\in\mathbb{Z},t\in\mathbb{R}$, ensuring that the boundary conditions of the original string vibrating equation Eq. (1a) are met. Therefore, we obtain the solution of Eq. (1a):
$u(x,t)=U(x,t),\quad\forall(x,t)\in[0,L]\times[0,T].$ (7)
Appendix E Benchmarks
We provide a detailed description of the three large-scale benchmarks in our experiments here: a challenging, self-generated heterogeneous 2D Navier-Stokes Equations dataset (HeterNS) and two large-scale benchmarks, one proposed by PDEformer (Ye et al., 2024), and the other collected by DPOT (Hao et al., 2024). These benchmarks cover a wide range of PDEs and diverse generalization scenarios, which can test the generalizability of PDE solvers well.
E.1 HeterNS
Similar to FNO (Li et al., 2021a), we consider the 2D Navier-Stokes equation in vorticity formulation for the viscous, incompressible fluid on a unit torus. We consider both in-distribution test and zero-shot generalization settings on HeterNS. See Figure 12 and 13 for a visual representation.
@@ E.1 HeterNS
The problem involves two key PDE components: the viscosity coefficient and the force term. We experiment with viscosity coefficients $\nu\in[8\times 10^{-6},2\times 10^{-3}]$ and force terms in the form $f(x)=0.1(\sin(\omega\pi(x_{1}+x_{2}))+\cos(\omega\pi(x_{1}+x_{2})))$. Specifically, our training set consists of $\nu\in\{1\times 10^{-5},5\times 10^{-5},1\times 10^{-4},5\times 10^{-4},1\times 10^{-3}\}$ and $\omega\in\{1,2,3\}$, resulting in 15 unique combinations of PDE components. For each combination, we generate 1000 samples, yielding a total of 15,000 training samples. The dataset can be accessed at the following anonymous link. [^2] https://drive.google.com/drive/folders/142c518gF9DWDD9FOx7TvtEwaNb5nHZUa
@@
E.2 1D Time-dependent PDEs
This benchmark is proposed by PDEformer (Ye et al., 2024). It contains 3 million high-quality 1D time-dependent PDEs with various equation components for training and then evaluates the model performance using in distribution test sets and zero-shot generalization performance on Burgers and Advection equation from PDEBench (Takamoto et al., 2022), which is another distinct benchmark. See Figure 14 for a visual representation.
@@
where $f_{i}(u)=c_{i1}u+c_{i2}u^{2}+c_{i3}u^{3},i=0,1$. Each coefficient $c_{ik}$ is set to zero with a probability of 0.5, and otherwise uniformly sampled from the interval $[-3,3]$. The variables $\kappa(x)$ and $s(x)$ can be zero, constant or physical fields, which are all randomly sampled from pre-defined distributions, as detailed in PDEformer’s original paper (Ye et al., 2024). The initial condition $g(x)$ is randomly generated within the family of trigonometric functions, a super-position of sinusoidal waves as, $u_{0}(x)=\sum_{k_{i}=k_{1},...,k_{N}}A_{i}\sin(k_{i}x+\phi_{i})$, where $k_{i}=2\pi{n_{i}}/L_{x}$ are wave numbers and ${n_{i}}\in\mathbb{N}$ are selected randomly in $[1,n_{\mathrm{max}}]$, which is same as the zero-shot generalization tasks from PDEBench (Takamoto et al., 2022).
@@
We employ the following two 1D PDE datasets from PDEBench (Takamoto et al., 2022) as zero-shot generalization tasks. All zero-shot generalization tasks follow periodic boundary conditions and the same initial condition family as the training dataset. The resolution of these samples is $1024\times 201$. For each dataset, we use 1000 test samples. We downsample the spatial resolution of these datasets to 256 and maintain the temporal resolution unchanged. The zero-shot PDEs consist of the Burgers equation and the Advection equation.
@@
We also provide fine-tuning results on 1D time-dependent PDEs in Appendix A. Compared to zero-shot generalization, we fine-tune the model using an additional 9,000 samples while testing on the same 1,000 samples.
@@
E.3 2D Mixed PDEs
This benchmark is collected by DPOT (Hao et al., 2024), which consists of the following 12 diverse subsets from 4 benchmarks. We only conduct in-distribution tests in the 2D mixed PDEs. Notably, the in- distribution test set also involves challenging variations in the PDE components. See Figure 15 for a visual representations.
FNO-$\nu$ (Li et al., 2021a)
@@
PDEBench (Takamoto et al., 2022)
The following three subsets are derived from PDEBench (Takamoto et al., 2022), encompassing three distinct equations: the compressible Navier-Stokes equation (CNS), the diffusion-reaction equation (DR), and the shallow-water equation (SWE). All datasets considered in PDEBench adhere to periodic boundary conditions. The spatial resolution of this benchmark is $128\times 128$.
@@
$\partial_{t}h+\nabla\cdot(h\bm{u})$ $=0,$ (12a)
$\partial_{t}(h\bm{u})+\nabla\cdot\left(\cfrac{1}{2}h\bm{u}^{2}+\cfrac{1}{2}g_{r}h^{2}\right)$ $=-g_{r}h\nabla b.$ (12b)
where $h$ describes the water depth, $b$ describes a spatially varying bathymetry, $g_{r}$ describes the gravitational acceleration, and $\nabla\cdot(h\bm{u})$ can be interpreted as the directional momentum. A key characteristic of this dataset is its long prediction horizon. The task of interest is to predict the future 91 timesteps of water depth based on the first 10 timesteps of observations. In in-distribution tests, the initial conditions vary across samples.
@@
$\partial_{t}\bm{v}$ $=-\bm{v}\cdot\nabla\bm{v}+\mu\nabla^{2}\bm{v}-\nabla p+\bm{f},$ (15a)
$\nabla\cdot\bm{v}$ $=0.$ (15b)
@@
$\partial_{t}(\rho\bm{u})+\nabla\cdot(\rho\bm{u}^{2})$ $=-\nabla p+\nabla\cdot\mu(\nabla\bm{u}+\nabla\bm{u}^{T}),$ (16a)
$\nabla\cdot(\rho\bm{u})$ $=0.$ (16b)
where $\rho$ is the constant density, $\mu$ is the dynamic viscosity, $\bm{u}=(u,v)^{T}$ is the velocity field, and $p$ is the pressure.
In in-distribution test settings, flows are generated for each problem with different PDE components, which are a combination of three types: (1) boundary conditions, (2) fluid physical coefficients such as density and viscosity, and (3) the geometry of the field. The boundary conditions refer to the inlet velocity or movement velocity, depending on the specific case. Each type of PDE component corresponds to a distinct subset. In each subset, the corresponding PDE components are varied while other parameters remain constant. We mix the three subsets following DPOT’s configuration (Hao et al., 2024), resulting in 9,000 training samples and 1,000 testing samples. The initial resolution is $64\times 64$, which is then interpolated to $128\times 128$. The task is to predict the next 10 timesteps of velocity given the first 10 timesteps of observations.
@@
We also provide fine-tuning results on 2D mixed PDEs in Appendix A. Given the significant diversity across the 12 subsets, we fine-tune the model using a specific training subset to allow it to focus on the target subset and achieve improved performance.
Appendix F Implementation Details

### v4

@@ Appendix D Full Results for the HeterNS Benchmark
Appendix E More Showcases
We provide additional showcases here to supplement the numerical results in main text. First, we visualize the in-distribution test and zero-shot generalization cases on HeterNS in Figure 13 and Figure 14 respectively. Next, we present visualizations for 1D time-dependent PDEs in Figure 15. Finally, we illustrate the 12 diverse datasets from 2D mixed PDEs in Figure 16.
Figure 13: Error maps (absolute difference between model predictions and ground truth) for in-distribution tests with top three baselines on HeterNS. See Table 10 and 11 for numerical comparison (relative L2). All data has the same initial condition and differs in viscosity ($\nu$) and force ($\omega$) (shown in the first row by the pairs ($\nu,\omega$)). Unisolver achieved the best visual performance among the compared baselines.
Figure 14: Error maps for zero-shot generalization settings with top three baselines on the HeterNS dataset with the same initial conditions and differs in viscosity ($\nu$) and force ($\omega$) (shown in the first row by the pairs ($\nu,\omega$)). See Table 10 and 11 for numerical comparison.
Figure 15: Error maps on the in-distribution test and zero-shot generalization (Burgers and Advection equation from PDEBench [37]) settings in 1D time-dependent PDEs. See Table 4 for numerical comparison. We visualize two cases: periodic boundary conditions and Robin boundary conditions in in-distribution tests. The number in the Burgers columns is the diffusion coefficient $\nu$ while the number in the Advection column is the advection speed $\beta$.
Figure 16: Unisolver predictions and error maps on 2D mixed PDEs. See Table 5 for numerical results. Both predictions and error maps are provided. As shown in the CFDBench-NS columns, Unisolver presents an impressive ability to handle different geometry conditions.
Appendix F Analytical Solution for the String Vibration Equation
$u(x,t)=\frac{1}{2}(\underbrace{\Phi(x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t)+\Phi(x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t)}_{\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}Initial position}}})+\frac{1}{2\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}}\underbrace{\int^{x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t}_{x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t}\Psi(\xi)}_{\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}Initial velocity}}}\mathrm{d}\xi+\frac{1}{2\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}}\underbrace{\int_{0}^{t}\mathrm{d}\tau\int_{x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}(t-\tau)}^{x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}(t-\tau)}}_{\text{{\color[rgb]{0.45,0.45,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.45,0.45,0.9}Geometry}}}\underbrace{f(\xi,\tau)}_{\text{{\color[rgb]{0.45,0.45,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.45,0.45,0.9}Force}}}\mathrm{d}\xi,\$ (3)
Detailed proof can be found in [8] or other relevant books.
Appendix G Benchmarks
We provide a detailed description of the three large-scale benchmarks in our experiments here: a challenging, self-generated heterogeneous 2D Navier-Stokes Equations dataset (HeterNS) and two large-scale benchmarks, one proposed by PDEformer [45], and the other collected by DPOT [11]. These benchmarks cover a wide range of PDEs and diverse generalization scenarios, which can test the generalizability of PDE solvers well.
G.1 HeterNS
Similar to FNO [19], we consider the 2D Navier-Stokes equation in vorticity formulation for the viscous, incompressible fluid on a unit torus. We consider both in-distribution test and zero-shot generalization settings on HeterNS. See Figure 13 and 14 for a visual representation.
@@ G.1 HeterNS
The problem involves two key PDE components: the viscosity coefficient and the force term. We experiment with viscosity coefficients $\nu\in[8\times 10^{-6},2\times 10^{-3}]$ and force terms in the form $f(x)=0.1(\sin(\omega\pi(x_{1}+x_{2}))+\cos(\omega\pi(x_{1}+x_{2})))$. Specifically, our training set consists of $\nu\in\{1\times 10^{-5},5\times 10^{-5},1\times 10^{-4},5\times 10^{-4},1\times 10^{-3}\}$ and $\omega\in\{1,2,3\}$, resulting in 15 unique combinations of PDE components. For each combination, we generate 1000 samples, yielding a total of 15,000 training samples. The dataset can be accessed at the following anonymous link. [^2] https://drive.google.com/drive/folders/1te5IyQHTznu_Kw7v3zDHg0i_KCHysPKw?usp=share_link
@@
G.2 1D Time-dependent PDEs
This benchmark is proposed by PDEformer [45]. It contains 3 million high-quality 1D time-dependent PDEs with various equation components for training and then evaluates the model performance using in distribution test sets and zero-shot generalization performance on Burgers and Advection equation from PDEBench [37], which is another distinct benchmark. See Figure 15 for a visual representation.
@@
where $f_{i}(u)=c_{i1}u+c_{i2}u^{2}+c_{i3}u^{3},i=0,1$. Each coefficient $c_{ik}$ is set to zero with a probability of 0.5, and otherwise uniformly sampled from the interval $[-3,3]$. The variables $\kappa(x)$ and $s(x)$ can be zero, constant or physical fields, which are all randomly sampled from pre-defined distributions, as detailed in PDEformer’s original paper [45]. The initial condition $g(x)$ is randomly generated within the family of trigonometric functions, a super-position of sinusoidal waves as, $u_{0}(x)=\sum_{k_{i}=k_{1},...,k_{N}}A_{i}\sin(k_{i}x+\phi_{i})$, where $k_{i}=2\pi{n_{i}}/L_{x}$ are wave numbers and ${n_{i}}\in\mathbb{N}$ are selected randomly in $[1,n_{\mathrm{max}}]$, which is same as the zero-shot generalization tasks from PDEBench [37].
@@
We employ the following two 1D PDE datasets from PDEBench [37] as zero-shot generalization tasks. All zero-shot generalization tasks follow periodic boundary conditions and the same initial condition family as the training dataset. The resolution of these samples is $1024\times 201$. For each dataset, we use 1000 test samples. We downsample the spatial resolution of these datasets to 256 and maintain the temporal resolution unchanged. The zero-shot PDEs consist of the Burgers equation and the Advection equation.
@@
We also provide fine-tuning results on 1D time-dependent PDEs in Appendix B. Compared to zero-shot generalization, we fine-tune the model using an additional 9,000 samples while testing on the same 1,000 samples.
@@
G.3 2D Mixed PDEs
This benchmark is collected by DPOT [11], which consists of the following 12 diverse subsets from 4 benchmarks. We only conduct in-distribution tests in the 2D mixed PDEs. Notably, the in- distribution test set also involves challenging variations in the PDE components. See Figure 16 for a visual representations.
FNO-$\nu$ [19]
@@
PDEBench [37]
The following three subsets are derived from PDEBench [37], encompassing three distinct equations: the compressible Navier-Stokes equation (CNS), the diffusion-reaction equation (DR), and the shallow-water equation (SWE). All datasets considered in PDEBench adhere to periodic boundary conditions. The spatial resolution of this benchmark is $128\times 128$.
@@
$\partial_{t}h+\nabla\cdot(h\boldsymbol{u})$ $=0,$ (8a)
$\partial_{t}(h\boldsymbol{u})+\nabla\cdot\left(\cfrac{1}{2}h\boldsymbol{u}^{2}+\cfrac{1}{2}g_{r}h^{2}\right)$ $=-g_{r}h\nabla b.$ (8b)
where $h$ describes the water depth, $b$ describes a spatially varying bathymetry, $g_{r}$ describes the gravitational acceleration, and $\nabla\cdot(h\boldsymbol{u})$ can be interpreted as the directional momentum. A key characteristic of this dataset is its long prediction horizon. The task of interest is to predict the future 91 timesteps of water depth based on the first 10 timesteps of observations. In in-distribution tests, the initial conditions vary across samples.
@@
$\partial_{t}\boldsymbol{v}$ $=-\boldsymbol{v}\cdot\nabla\boldsymbol{v}+\mu\nabla^{2}\boldsymbol{v}-\nabla p+\boldsymbol{f},$ (11a)
$\nabla\cdot\boldsymbol{v}$ $=0.$ (11b)
@@
$\partial_{t}(\rho\boldsymbol{u})+\nabla\cdot(\rho\boldsymbol{u}^{2})$ $=-\nabla p+\nabla\cdot\mu(\nabla\boldsymbol{u}+\nabla\boldsymbol{u}^{T}),$ (12a)
$\nabla\cdot(\rho\boldsymbol{u})$ $=0.$ (12b)
where $\rho$ is the constant density, $\mu$ is the dynamic viscosity, $\boldsymbol{u}=(u,v)^{T}$ is the velocity field, and $p$ is the pressure.
In in-distribution test settings, flows are generated for each problem with different PDE components, which are a combination of three types: (1) boundary conditions, (2) fluid physical coefficients such as density and viscosity, and (3) the geometry of the field. The boundary conditions refer to the inlet velocity or movement velocity, depending on the specific case. Each type of PDE component corresponds to a distinct subset. In each subset, the corresponding PDE components are varied while other parameters remain constant. We mix the three subsets following DPOT’s configuration [11], resulting in 9,000 training samples and 1,000 testing samples. The initial resolution is $64\times 64$, which is then interpolated to $128\times 128$. The task is to predict the next 10 timesteps of velocity given the first 10 timesteps of observations.
@@
We also provide fine-tuning results on 2D mixed PDEs in Appendix B. Given the significant diversity across the 12 subsets, we fine-tune the model using a specific training subset to allow it to focus on the target subset and achieve improved performance.
Appendix H Implementation Details